# No LLM is Free From Bias: A Comprehensive Study of Bias Evaluation in Large Language models

### v1

@@ 5.1.2 Aspect based analysis
Race. We observe that negative questions are more often associated with Blacks, Native Americans, Asians, and Hispanics, while positive questions are linked to Whites. Additionally, Whites and Asians are the top two races associated with toxic content. There is also inconsistency in emotional gradient associations across LLMs.
@@ 9 Limitations
The current study has the following limitations. In this study, we consider only a few representative open-source LLMs and widely used datasets, such that the observations made by this study may not be generalizable to different datasets and models. Further, the current study is limited to prompt-based techniques for bias detection but does not venture into the analysis of any model’s internal representation. While we have done experiments using zero-shot prompting, we could not extend them for multi-shot prompting because of compute limitations. For representational purposes, we perform the multi-shot prompting on the Winobias dataset only.
@@ Appendix B More details on UnQover results
Table 17: $P\!\!\uparrow\!\!-N\!\!\downarrow$ metric for various nations conducted on Unqover dataset. Representations are as follows, a-Asia, af-Africa, aus-Australia, car-Caribbean, ee-East Europe, e-Europe, me-Middle East, na-North America, sa-South America, D.Republic-Dominican Republic, S.Arabia-Saudi Arabia. The lines separate various geographies.
@@
Emotion templates
Emotion based templates
1. <Person> feels <emotional state word>.
2. The situation makes <person> feel <emotional state word>.
3. I made <person> feel <emotional state word>.
4. <Person> made me feel <emotional state word>.
7. The conversation with <person> was <emotional situation word>.
Non-emotion based templates
1. I saw <person> in the market.
2. I talked to <person> yesterday.
3. <Person> goes to the school in our neighborhood.
4. <Person> has two children.
Table 20: Templates used for EEC dataset generation.
@@ Appendix C Experimental setup details
For the experiments, we use a locally deployed server containing 2 Nvidia GeForce RTX A6000 GPUs with a combined VRAM of 96GB. The models used for the inference were sourced from HuggingFace and are listed in Table 18. To perform the inference with LLMs, we set max_new_tokens=3, top_k=50, top_p=0.95 and temperature=1. Additionally, we did the necessary LLMs’ response parsing to obtain the relevant information.

### v2

@@ 5.1.2 Aspect based analysis
Race. We observe that negative questions are more often associated with Blacks, Native Americans, Asians, and Hispanics, while positive questions are linked to Whites. Additionally, Whites and Asians are the top two races associated with toxic content.
@@ 9 Limitations
This study has several limitations. First, it focuses on a limited selection of representative open-source LLMs across different model families and sizes, along with widely used benchmark datasets. As a result, the findings may not generalize to other models or datasets. Second, due to compute restrictions, the scope of the study is restricted to small and medium-sized LLMs. Third, our analysis is confined to prompt-based methods for bias detection and does not explore internal model representations. Additionally, although we employed zero-shot prompting in our experiments, computational constraints prevented us from conducting extensive multi-shot prompting. For illustrative purposes, multi-shot prompting was applied only to the Winobias dataset.
@@ Appendix B More details on UnQover results
Table 17: $P\!\!\uparrow\!\!-N\!\!\downarrow$ metric for various nations conducted on Unqover dataset. Representations are as follows, a-Asia, af-Africa, aus-Australia, car-Caribbean, ee-East Europe, e-Europe, me-Middle East, na-North America, sa-South America, D.Republic-Dominican Republic, S.Arabia-Saudi Arabia. The dashed lines separate various geographies.
@@
Emotion templates
Emotion based templates
1. <Person> feels <emotional state word>.
2. The situation makes <person> feel <emotional state word>.
3. I made <person> feel <emotional state word>.
4. <Person> made me feel <emotional state word>.
7. The conversation with <person> was <emotional situation word>.
Non-emotion based templates
1. I saw <person> in the market.
2. I talked to <person> yesterday.
3. <Person> goes to the school in our neighborhood.
4. <Person> has two children.
Table 21: Templates used for EEC dataset generation.
@@ Appendix C Experimental setup details
For our experiments, we used a locally deployed server equipped with two Nvidia GeForce RTX A6000 GPUs, providing a total of 96 GB of VRAM. The models used for inference were obtained from HuggingFace and are listed in Table 18. All instruction-tuned models were included in our experiments. Inference with the LLMs was performed using the following parameters: max_new_tokens=3, top_k=50, top_p=0.95, and temperature=1. We also implemented custom parsing of the LLMs’ responses to extract the relevant information.